# Fisher Mask Nodes for Language Model Merging

### 14 Mar 2024 v1

@@ Abstract
Fine-tuning pre-trained models provides significant advantages in downstream performance. The ubiquitous nature of pre-trained models such as BERT and its derivatives in natural language processing has also led to a proliferation of task-specific fine-tuned models. As these models typically only perform one task well, additional training or ensembling is required in multi-task scenarios. The growing field of model merging provides a solution, dealing with the challenge of combining multiple task-specific models into a single multi-task model. In this study, we introduce a novel model merging method for Transformers, combining insights from previous work in Fisher-weighted averaging and the use of Fisher information in model pruning. Utilizing the Fisher information of mask nodes within the Transformer architecture, we devise a computationally efficient weighted-averaging scheme. Our method exhibits a regular and significant performance increase across various models in the BERT family, outperforming full-scale Fisher-weighted averaging in a fraction of the computational cost, with baseline performance improvements of up to +6.5 and a speedup of 57.4x. Our results prove the potential of our method in current multi-task learning environments and suggest its scalability and adaptability to new model architectures and learning scenarios.
@@ 5.   Results
We note that our method also provides a significant efficiency advantage over Fisher-weighted averaging, with a 57.4x mean speedup in the case of BERT (Large). To reduce the effect of overhead, the speedup is calculated using only the average runtime of the Fisher information approximation algorithm for each method, with an equal number of samples, running on a Nvidia T4 GPU.

### 18 Mar 2024 v2

@@ Abstract
Fine-tuning pre-trained models provides significant advantages in downstream performance. The ubiquitous nature of pre-trained models such as BERT and its derivatives in natural language processing has also led to a proliferation of task-specific fine-tuned models. As these models typically only perform one task well, additional training or ensembling is required in multi-task scenarios. The growing field of model merging provides a solution, dealing with the challenge of combining multiple task-specific models into a single multi-task model. In this study, we introduce a novel model merging method for Transformers, combining insights from previous work in Fisher-weighted averaging and the use of Fisher information in model pruning. Utilizing the Fisher information of mask nodes within the Transformer architecture, we devise a computationally efficient weighted-averaging scheme. Our method exhibits a regular and significant performance increase across various models in the BERT family, outperforming full-scale Fisher-weighted averaging in a fraction of the computational cost, with baseline performance improvements of up to +6.5 and a speedup of 57.4x in the biggest model. Our results prove the potential of our method in current multi-task learning environments and suggest its scalability and adaptability to new model architectures and learning scenarios.
@@ 1.   Introduction
Our code is publicly released on Github. [^1] https://github.com/thennal10/fisher-nodes-merging
@@ 5.   Results
We note that our method also provides a significant efficiency advantage over Fisher-weighted averaging, with a 57.4x elapsed time speed-up for BERT (Large). To reduce the effect of overhead, the speedup is calculated using only the average runtime of the Fisher information approximation algorithm for each method, with an equal number of samples, running on a Nvidia T4 GPU.
@@ 8.   Language Resource References
Appendix
1.   Checkpoints
For all architectures we mention, 6 models finetuned on the 6 respective tasks from the GLUE dataset were used. The checkpoints are given below.
1.1.   BERT Base
  • MNLI: JeremiahZ/bert-base-uncased-mnli
  • QQP: textattack/bert-base-uncased-QQP
  • QNLI: textattack/bert-base-uncased-QNLI
  • SST2: textattack/bert-base-uncased-SST-2
  • MRPC: textattack/bert-base-uncased-MRPC
  • RTE: textattack/bert-base-uncased-RTE
1.2.   BERT Large
  • MNLI: yoshitomo-matsubara/bert-large-uncased-mnli
  • QQP: yoshitomo-matsubara/bert-large-uncased-qqp
  • QNLI: yoshitomo-matsubara/bert-large-uncased-qnli
  • SST2: yoshitomo-matsubara/bert-large-uncased-sst2
  • MRPC: yoshitomo-matsubara/bert-large-uncased-mrpc
  • RTE: yoshitomo-matsubara/bert-large-uncased-rte
1.3.   BERT Tiny
  • MNLI: M-FAC/bert-tiny-finetuned-mnli
  • QQP: M-FAC/bert-tiny-finetuned-qqp
  • QNLI: M-FAC/bert-tiny-finetuned-qnli
  • SST2: M-FAC/bert-tiny-finetuned-sst2
  • MRPC: M-FAC/bert-tiny-finetuned-mrpc
  • RTE: muhtasham/bert-tiny-finetuned-glue-rte
1.4.   RoBERTa
  • MNLI: JeremiahZ/roberta-base-mnli
  • QQP: JeremiahZ/roberta-base-qqp
  • QNLI: JeremiahZ/roberta-base-qnli
  • SST2: JeremiahZ/roberta-base-sst2
  • MRPC: JeremiahZ/roberta-base-mrpc
  • RTE: JeremiahZ/roberta-base-rte